# SimWorld-Robotics: Synthesizing Photorealistic and Dynamic Urban Environments for Multimodal Robot Navigation and Collaboration

### v1

@@ A.8.3 New Tasks
Defining tasks in SWR is similar to standard Gym environments [brockman2016openaigym], as the Python API of SWR follows the same format for agent control (including pedestrians and vehicles). Users can spawn different types of agents, customize observation spaces (e.g., RGB, depth, or semantic segmentation) and action spaces (continuous or discrete actions), and program new goal definitions (e.g., language instructions, target images, or spatial objectives). Additionally, while the current pedestrians follow rule-based logic, users can override behaviors to simulate complex or rare cases. For instance, one can simulate a jaywalking pedestrian by scripting a few agents to cross during a red light, while placing a robot at the intersection to evaluate its reaction.
@@ A.9 SWR Technical and Operational Details
Python API and runtime environment. On top of UE5, we’ve implemented a dedicated Python layer that communicates with the engine through an updated UnrealCV‑based TCP server [10.1145/3123266.3129396], where high‑level commands issued in Python are forwarded over TCP straight into the UE5 runtime. On top of this API, we provide a standard Gym interface so that researchers can plug in and benchmark any baseline with minimal effort. We will distribute a Windows executable and a Singularity container for Ubuntu and macOS so that users can run SWR out of the box without local compilation; all binaries, container images, and the Gym wrapper will be open‑sourced upon publication.
@@ B.1 Comparing SimWorld-MMNav with Prior Vision-Language Navigation Benchmarks
CVDN [thomason2019visionanddialognavigation] Indoor, Static 90 Camera Graph-based Manual 7 7,415
REVERIE[qi2020reverieremoteembodiedvisual] Indoor, Static 90 Camera Graph-based Manual 5 7,000
TouchDown [chen2020touchdownnaturallanguagenavigation] Outdoor, Static 1 Camera Graph-based Procedural 35 9,326
ANDH [fan2023aerialvisionanddialognavigation] Outdoor, Static 1 Drone 6 Manual 7 6,269
AerialVLN [liu2023aerialvlnvisionandlanguagenavigationuavs] Outdoor, Static 25 Drone 8 Manual 204 8,446
@@ B.2 Comparing SimWorld-MRS with Prior Multi-Robot Collaboration and Navigation Benchmarks
RoCo [beck2023roco] Indoor, Static Room 2 Mobile arm None None Collab. pick&place
DOROTHIE [ma2022dorothie] Outdoor, Dynamic Road 1 (ego car) Vehicle Spoken dlg. $\sim\!20$ tok. Full Dialogue nav.
MetaUrban [wu2024metaurban] Outdoor, Dynamic City 1 + E-scooter/robot None Full Micromobility nav.
DriVLMe [driVLMe2023] Outdoor, Dynamic City $10+$ Vehicle None Full Coop. driving
RobotSlang [kapoor2024robotslang] Indoor, Static Lab 2–3 Mobile base Natural-lang. 80 char Partial Joint object search
Where Are You? [das2018where] Outdoor, Static City 2 (tourist+guide) Pedestrian Natural-lang. 40 tok. Split Localization via dialog
@@ C.5 Baseline Detail
To reduce the mapping complexity and prevent hallucinations, we decompose step-wise action prediction in multimodal instruction-following into two modules: a context-aware perception module and a ReAct-based [yao2023reactsynergizingreasoningacting] action module as detailed in Figure 10. Throughout this process, the model autonomously maintains a working memory.
@@
During finetuning, the model receives as input the current observation, the target image representing the expected view upon subtask completion, the instruction, the current orientation, and the ground-truth action history. Based on these inputs, the model is trained to predict the estimated distance, the anticipated final orientation, and the remaining action sequence as a form of CoT planning, followed by the prediction of the next action. We finetune the Qwen2.5-VL-7B-Instruct model using LoRA [dettmers2023qloraefficientfinetuningquantized] applied to both the language model head and the merging projection layer. The loss is computed solely on the tokens generated by the decoder. We used four A100 GPUs with 80GB VRAM each for finetuning. The hyperparameters can be found in Table 10
@@ C.6 More Quantitative Results
We also provide the 95% CI of our main results, as Table 12 shows. For success rate and subtask success rate, we use binomial proportion confidence interval [wilson-ci] and display both lower and upper bound in the table.
@@ D.4 Baseline Detail
The ROCO-based [beck2023roco] setting extends the oracle setup by introducing collaborative planning and communication between two robots. After the two agents communicate and confirm the follower’s current location in terms of a landmark building, the main robot uses access to the ground-truth map to identify the landmark closest to the midpoint between both agents. It then performs path planning to compute the optimal trajectories for both agents to converge at this intermediate landmark. The main robot transforms the computed path for the follower into natural language instructions, which are communicated to the follower robot. The follower then performs multimodal instruction following to navigate toward the meeting point and rendezvous with the main robot. An example for communication in the ROCO baseline can be seen in Figure 11.

### v2

@@ A.8.3 New Tasks
Defining tasks in SWR is similar to standard Gym environments [7], as the Python API of SWR follows the same format for agent control (including pedestrians and vehicles). Users can spawn different types of agents, customize observation spaces (e.g., RGB, depth, or semantic segmentation) and action spaces (continuous or discrete actions), and program new goal definitions (e.g., language instructions, target images, or spatial objectives). Additionally, while the current pedestrians follow rule-based logic, users can override behaviors to simulate complex or rare cases. For instance, one can simulate a jaywalking pedestrian by scripting a few agents to cross during a red light, while placing a robot at the intersection to evaluate its reaction.
@@ A.9 SWR Technical and Operational Details
Python API and runtime environment. On top of UE5, we’ve implemented a dedicated Python layer that communicates with the engine through an updated UnrealCV‑based TCP server [41], where high‑level commands issued in Python are forwarded over TCP straight into the UE5 runtime. On top of this API, we provide a standard Gym interface so that researchers can plug in and benchmark any baseline with minimal effort. We will distribute a Windows executable and a Singularity container for Ubuntu and macOS so that users can run SWR out of the box without local compilation; all binaries, container images, and the Gym wrapper will be open‑sourced upon publication.
@@ B.1 Comparing SimWorld-MMNav with Prior Vision-Language Navigation Benchmarks
CVDN [49] Indoor, Static 90 Camera Graph-based Manual 7 7,415
REVERIE[40] Indoor, Static 90 Camera Graph-based Manual 5 7,000
TouchDown [9] Outdoor, Static 1 Camera Graph-based Procedural 35 9,326
ANDH [15] Outdoor, Static 1 Drone 6 Manual 7 6,269
AerialVLN [30] Outdoor, Static 25 Drone 8 Manual 204 8,446
@@ B.2 Comparing SimWorld-MRS with Prior Multi-Robot Collaboration and Navigation Benchmarks
RoCo [59] Indoor, Static Room 2 Mobile arm None None Collab. pick&place
DOROTHIE [32] Outdoor, Dynamic Road 1 (ego car) Vehicle Spoken dlg. $\sim\!20$ tok. Full Dialogue nav.
MetaUrban [54] Outdoor, Dynamic City 1 + E-scooter/robot None Full Micromobility nav.
DriVLMe [21] Outdoor, Dynamic City $10+$ Vehicle None Full Coop. driving
RobotSlang [5] Indoor, Static Lab 2–3 Mobile base Natural-lang. 80 char Partial Joint object search
Where Are You? [11] Outdoor, Static City 2 (tourist+guide) Pedestrian Natural-lang. 40 tok. Split Localization via dialog
@@ C.5 Baseline Detail
To reduce the mapping complexity and prevent hallucinations, we decompose step-wise action prediction in multimodal instruction-following into two modules: a context-aware perception module and a ReAct-based [56] action module as detailed in Figure 10. Throughout this process, the model autonomously maintains a working memory.
@@
During finetuning, the model receives as input the current observation, the target image representing the expected view upon subtask completion, the instruction, the current orientation, and the ground-truth action history. Based on these inputs, the model is trained to predict the estimated distance, the anticipated final orientation, and the remaining action sequence as a form of CoT planning, followed by the prediction of the next action. We finetune the Qwen2.5-VL-7B-Instruct model using LoRA [13] applied to both the language model head and the merging projection layer. The loss is computed solely on the tokens generated by the decoder. We used four A100 GPUs with 80GB VRAM each for finetuning. The hyperparameters can be found in Table 10
@@ C.6 More Quantitative Results
We also provide the 95% CI of our main results, as Table 12 shows. For success rate and subtask success rate, we use binomial proportion confidence interval [53] and display both lower and upper bound in the table.
@@ D.4 Baseline Detail
The ROCO-based [59] setting extends the oracle setup by introducing collaborative planning and communication between two robots. After the two agents communicate and confirm the follower’s current location in terms of a landmark building, the main robot uses access to the ground-truth map to identify the landmark closest to the midpoint between both agents. It then performs path planning to compute the optimal trajectories for both agents to converge at this intermediate landmark. The main robot transforms the computed path for the follower into natural language instructions, which are communicated to the follower robot. The follower then performs multimodal instruction following to navigate toward the meeting point and rendezvous with the main robot. An example for communication in the ROCO baseline can be seen in Figure 11.